OpenStudy (anonymous):

John is 12 years older than Jane. In 2 years, John's age will be four times larger than what Jane's age was 4 years ago. How old is John now?

OpenStudy (anonymous):

let,jane age=x john age=x+12 after 2 yrs... x+14=4(x-4) x-4x=-16-14 -3x=-30 x=10 john age=22

OpenStudy (anonymous):

Thank you!! I hate word problems :p